# Distributionally Robust Acceleration Control Barrier Filter for Efficient UAV Obstacle Avoidance

### v1

@@ Abstract
Dynamic obstacle avoidance (DOA) for unmanned aerial vehicles (UAVs) requires fast reaction under limited onboard resources. We introduce the distributionally robust acceleration control barrier function (DR-ACBF) as an efficient collision avoidance method maintaining safety regions. The method constructs a second-order control barrier function as linear half-space constraints on commanded acceleration. Latency, actuator limits, and obstacle accelerations are handled through an effective clearance that considers dynamics and delay. Uncertainty is mitigated using Cantelli tightening with per-obstacle risk. A DR-conditional value at risk (DR-CVaR)- based early trigger expands margins near violations to improve DOA. Real-time execution is ensured via constant-time Gauss-Southwell projections. Simulation studies achieve similar avoidance performance at substantially lower computational effort than state-of-the-art baseline approaches. Experiments with Crazyflie drones demonstrate the feasibility of our approach.
@@ I Introduction
Unmanned aerial vehicles (UAVs) are increasingly deployed in day-to-day services that require rapid responses to dynamic obstacles. Applications include urban UAV logistics, disaster response, and inspection in cluttered environments where obstacles such as other drones, birds, or humans may appear or move unpredictably [Rinner_Computer2021]. To operate safely in these cases, UAVs must react fast to avoid collisions while maintaining stability. Dynamic obstacle avoidance (DOA) relies on high-rate perception and control and must handle perception uncertainty, latency, and limited computation under tight real-time constraints. Even a small perception-to-actuation latency can cause severe safety violations at high speeds [falanga2019ral]. In addition, vision and radio detection and ranging (RADAR) sensors produce high-rate measurements that further strain the capabilities of current DOA systems [mandaokar_widroit2025, falanga2019ral].
@@
Classical UAV sense-and-avoid approaches rely on trajectory optimization [chen2019robio, foam2021, rota_icra2024], nonlinear model predictive control (NMPC) [lindqvist2020ral], or reactive planners [lin2019fast3d, singletary2021iros]. While these methods can handle static or slowly moving obstacles, they are computationally demanding or rely on conservative safety margins, making them unsuitable for agile flights. Moreover, these optimization-based approaches often lack formal safety guarantees in the presence of fast and uncertain obstacle motion. Recently, control barrier functions (CBFs) emerged as a reliable framework to guarantee safety by ensuring forward invariance of a safe set [AdaptiveCBF2022, ames2017tac, ames2014acc] throughout the entire mission. The classical CBF is implemented using a quadratic program (CBF-QP) and has been applied to ground robots, manipulators, and UAVs to enforce collision avoidance in real-time. Similarly, the existence of higher-order CBF (HOCBF) extends classical CBFs to handle higher relative-degree systems [AdaptiveCBF2022], which is crucial for UAVs with second-order dynamics. The implemented velocity-based formulations, such as VOCBF [VOCBF2025], encode safety in velocity space. Moreover, the CBF is integrated with differentiable optimization-based methods to enable smoother control pipelines [dai2023ral]. Similarly, a probabilistic enumeration is introduced for stochastic risk allocation [ProbEnum2025]. Runtime adaptive safety filtering and robust formulations have further broadened the applicability of CBFs in robotics [SafetyFly2025, RobustCBF2025].
@@
Despite this progress, the current CBF implementation faces limitations for DOA in UAVs. Existing CBF formulations act at the position or velocity level [VOCBF2025, FixedWingCBF2025], but do not provide closed-form acceleration-space constraints compatible with fast low-level thrust control. Although HOCBF implies extension of classical CBF to higher-order systems, it does not target acceleration or thrust as a control variable for DOA in UAVs. The HOCBF ensures safety at higher relative degrees [AdaptiveCBF2022], but requires solving QPs online, which is unsuitable for embedded flight controllers. Current implementations of CBF insufficiently address latency and uncertainty for DOA in UAVs. The ISSf-CBF [ISSfCBF2023] and tube-based formulations [TubeCBF2023] consider robustness to input delay, but they are conservative and rely on iterative solvers. Probabilistic CBFs improve risk handling, but have not been adapted to acceleration-level constraints or integrated with hard real-time systems [ProbEnum2025, SafetyFly2025]. Similarly, CBF with chance-constraints shows the importance of probabilistic risk margins for UAVs but remains velocity-centric [cccbf]. The QP-based CBF solvers and sampling-based planners are too slow for high-rate DOA control, particularly on resource-constrained UAV hardware [FixedWingCBF2025, RobustCBF2025, AdaptiveCBF2022]. There are lighter and computationally efficient extensions of MPC, such as TinyMPC, which has been tested on Crazyflie but is under scrutiny for its fast DOA in UAVs [tinyMPC].
@@
In this paper, we address the current CBF-QP limitation for DOA in UAVs with a distributionally robust acceleration control barrier function (DR-ACBF). DR-ACBF directly restricts UAV acceleration commands through safe, linear, half-space constraints. The half-space constraints specify the safe side of the projection plane for the UAV’s acceleration, ensuring that the UAV remains within this region. We demonstrate that enforcing a linear acceleration-space constraint derived from the second-order CBF dynamics satisfies the original CBF inequality over a short horizon under bounded uncertainty and latency assumptions. Additionally, the perception-to-actuation latency and uncertainty are handled by incorporating clearance buffering [falanga2019ral] and probabilistic tightening based on Cantelli inequalities [Cantelli], similar in spirit to stochastic CBFs [ProbEnum2025, SafetyFly2025] but specialized for acceleration space. Moreover, an early-warning layer based on distributionally robust conditional value at risk (DR-CVaR) [cvar2023, drcvar2023, wasserstein] is integrated to anticipate and adjust safety margins before any violation becomes imminent. Instead of solving QPs, we implement a constant-time Gauss-Southwell projection [gauss, fastproj] to enforce constraints, enabling real-time onboard execution.
@@
In contrast to existing CBF-driven collision avoidance frameworks, our DR-ACBF framework provides: (i) explicit acceleration-space tightening to stay safe, (ii) low computational cost via cyclic projection, and (iii) probabilistic and latency-aware safety margins. We derive a closed-form line-of-sight (LoS) acceleration half-space with horizon-matched gains and a fixed-time projection enforcement with Gauss-Southwell. To the best of our knowledge, no prior works present an acceleration-space CBF filter with latency and risk tightening for UAV obstacle avoidance. The DR-ACBF achieves real-time, reliable UAV obstacle avoidance via acceleration-level CBFs, DR-CVaR trigger, and constant-time projection, reducing command time by over 37% while maintaining high safety and success rates as evaluated in a simulation study. Experiments with Crazyflie drones demonstrate the feasibility of our approach.
@@ II Problem Formulation
Figure 1: Illustration of the proposed DR-ACBF framework. The UAV avoids dynamic obstacles (top) by constraining its acceleration within distributionally robust half-spaces marked in grey. The UAV view (bottom) depicts the projected half-space plane towards the obstacles. The effective clearance $R_{\mathrm{eff}}$ accounts for latency and actuation limits, while the obstacles’ uncertainty ellipsoids $\sigma_{dH,i}$ and risk levels $\alpha_{i}$ tighten the constraints under sensing uncertainty. The safe set of action $C$ ensures that the UAV’s safe zone is not violated.
@@ II-A Problem Statement
We model the UAV dynamics by second-order differential equations for simpler control design. We consider an acceleration-controlled quadrotor UAV operating in an environment with $N$ dynamic obstacles. The obstacles are detected and tracked with typical trackers (e.g., SAFE-IMM tracker [mandaokar_safeimm]), which deliver state estimates of the obstacles. The input space consists of the position $p_{e}\in\mathbb{R}^{3}$, velocity $v_{e}\in\mathbb{R}^{3}$, and acceleration $a_{e}\in\mathbb{R}^{3}$ of the UAV as well as the relative position, velocity, and acceleration of the obstacle. A nominal controller provides an acceleration command $a_{\mathrm{n}}\in\mathbb{R}^{3}$, while the UAV is subject to perception-to-actuation latency $\Delta\!\geq\!0$ and bounded actuation, i.e., limits for acceleration $a_{\max}\!>\!0$, jerk $j_{\max}\!>\!0$, and speed $v_{\max}\!>\!0$. The objective of DR-ACBF is to design a real-time safety filter that modifies $a_{\mathrm{n}}$ to a safe command $a_{e}\in\mathbb{R}^{3}$ while enforcing probabilistic collision avoidance under a user-specified global risk budget $\alpha_{\mathrm{total}}$ (see Figure 1).
@@ II-D Classical CBF Formulation
Ames et al. [ames2014acc, ames2017tac] introduced the classical control Lyapunov functions (CLF)-CBF-QP architecture, which guarantees that an autonomous agent moves toward its goal while maintaining safety constraints. Let $x\in\mathbb{R}^{n}$ denote the system state and $u\in\mathbb{R}^{m}$ the control input. The control-affine dynamics are
@@
When (9) holds, the safe set $\mathcal{C}$ is forward invariant. If $x(0)\in\mathcal{C}$, then $x(t)\in\mathcal{C}$ for all $t\geq 0$. A continuously differentiable $V:\mathbb{R}^{n}\!\to\!\mathbb{R}_{\geq 0}$ measures the distance to the desired goal state $x_{\mathrm{des}}$. Therefore, the Lie derivatives are $L_{f}V(x)=\nabla V(x)^{\top}f(x)$ and $L_{g}V(x)=\nabla V(x)^{\top}g(x)$. $V$ is a CLF with a Lyapunov decay rate $c>0$ such that
$\inf_{u}\big[L_{f}V(x)+L_{g}V(x)u+cV(x)\big]\leq 0.$ (10)
This constraint ensures that some admissible $u$ yields $\dot{V}(x)\leq-cV(x)$ and thus stabilizes the system. At each control step, one computes $u^{\star}$ by solving the QP
$(u^{\star},\delta^{\star})=\arg\min_{u,\delta}\tfrac{1}{2}\|u-u_{\mathrm{nom}}\|^{2}+W\,\delta^{2}$ (11)
$\text{s.t.}\quad L_{f}V(x)+L_{g}V(x)u+cV(x)\leq\delta,$
$L_{f}h(x)+L_{g}h(x)u+\alpha(h(x))\geq 0,$
$\delta\geq 0.$
where $u_{\mathrm{nom}}$ is a nominal control, $W$ is a weighting coefficient, and $\delta$ relaxes the goal constraint. The CBF inequality (9) strictly enforces safety while the CLF (10) is a soft constraint, i.e., performance is relaxed when safety requires it, as in (11).
@@ III Method
Figure 2: System block diagram. The process begins with sensing the UAV and obstacle states. UAV control generates a nominal acceleration command $a_{n}$, which is corrected by DR-ACBF avoidance. The DR-CVaR layer provides an early warning, the acceleration estimation, and risk allocation to compute the safety constraints. Finally, ACBF and projection replace $a_{n}$ with the safe acceleration command $a_{e}$ to avoid obstacles.
@@ III-A Acceleration CBF
We model the UAV dynamics similar to Equations (3) with $\|a_{e}\|\leq a_{\max}$:
@@
The combined sum of obstacle and UAV radius is $R_{\mathrm{sum}}=R_{\mathrm{uav}}+R_{\mathrm{obs}}$ as illustrated in Figure 1. Then the barrier function for distance is
@@
where $k_{1}$ damps the approach velocity and $k_{2}$ penalizes distance violation. In order to project the second-order condition into an acceleration half-space, we linearize it over a short lookahead horizon $H>0$ and eliminate the obstacle acceleration $a_{i}$ by worst-case padding, yielding an affine constraint in the UAV acceleration
@@
where $R_{\mathrm{eff},i}$ is the effective clearance that accounts for sensing and actuation latencies. Each obstacle thus defines one linear half-space of safe accelerations. The effective clearance in distance is enlarged to account for perception latency, actuation limits, and jerk bound:
@@
where $v_{\max}$ is the UAV speed limit, $a_{\max}$ and $j_{\max}$ are acceleration and jerk bounds, $\Delta$ is total perception-to-actuation latency, and the $\hat{a}_{\max,i}$ is the learned bound on obstacle acceleration. The sliding-mode differentiator (SMD) is used to compute the $\hat{a}_{\max,i}$ with a user-defined $a_{\max,i}$, aiming to keep it large to address the worst possible acceleration of obstacles. This is explained in detail in the next sections.
In order to compensate for the system’s latency changes, the latency contribution inside $R_{\mathrm{eff}}$ is bounded by $v_{\max}\Delta+\tfrac{1}{2}a_{\max}\Delta^{2}$. While the new command is held over $H$ with bounded jerk $j_{\max}$, the worst additional shrink from the acceleration ramp is upper-bounded by $\tfrac{1}{6}j_{\max}H^{3}$. The obstacle displacement over $H$ is limited by $\tfrac{1}{2}\hat{a}_{\max,i}H^{2}$. Summing up all these components to $R_{\mathrm{sum}}$ yields (18). The $\Delta$ includes pre-actuation latency to account for delays in activating the controls. The impact of such latency on safe high-speed navigation with bounded acceleration is analyzed in [falanga2019ral]. The trackers provide estimates of position and velocity [mandaokar_safeimm]. Therefore, if we consider the position and velocity covariance from the tracker, the LoS variance is bounded by Eq. (6).
Over a short lookahead horizon $H$, the coefficients $k_{1}=2/H$ and $k_{2}=2/H^{2}$ ensure that any acceleration satisfying the affine constraint maintains a non-decreasing barrier value of $h_{i}(t)$ over $H$. These coefficients follow from the second-order Taylor expansion of $h_{i}(t)$ over the horizon $H$. Enforcing the short-horizon condition $h_{i}(t+H)\geq 0$ yields $k_{1}$ and $k_{2}$.

### v2

@@ Abstract
Dynamic obstacle avoidance (DOA) for unmanned aerial vehicles (UAVs) requires fast reaction under limited onboard resources. We introduce the distributionally robust acceleration control barrier function (DR-ACBF) as an efficient collision avoidance method maintaining safety regions. The method constructs a second-order control barrier function as linear half-space constraints on commanded acceleration. Latency, actuator limits, and obstacle accelerations are handled through an effective clearance that considers dynamics and delay. Uncertainty is mitigated using Cantelli tightening with per-obstacle risk. A DR-conditional value at risk (DR-CVaR) early trigger expands margins near violations to improve DOA. To meet real-time avoidance-control at 100 Hz, we use fixed-time Gauss-Southwell projections instead of quadratic programs (QP). Simulation results show similar avoidance performance with 31% lower computational load than QP and outperform the state-of-the-art baseline approaches. Experiments with Crazyflie UAVs demonstrate the feasibility of our approach.
@@ I Introduction
Unmanned aerial vehicles (UAVs) are increasingly deployed in day-to-day services that require rapid responses to dynamic obstacles. Applications include urban UAV logistics, disaster response, and inspection in cluttered environments where obstacles such as other UAVs, birds, or humans may appear or move unpredictably [30]. To operate safely in these cases, UAVs must react fast to avoid collisions while maintaining stability. Dynamic obstacle avoidance (DOA) relies on high-rate perception and control and must handle perception uncertainty, latency, and limited computation under tight real-time constraints. Even a small perception-to-actuation latency can cause severe safety violations at high speeds [9]. In addition, vision and radio detection and ranging (RADAR) sensors produce high-rate measurements that further strain the capabilities of current DOA systems [17, 9].
@@
Classical UAV sense-and-avoid approaches rely on trajectory optimization [4, 10, 35], nonlinear model predictive control (NMPC) [16], or reactive planners [15, 32]. While these methods can handle static or slowly moving obstacles, they are computationally demanding or rely on conservative safety margins, making them unsuitable for agile flights. Moreover, these optimization-based approaches often lack formal safety guarantees in the presence of fast and uncertain obstacle motion. Recently, control barrier functions (CBFs) emerged as a reliable framework to guarantee safety by ensuring forward invariance of a safe set [34, 3, 2] throughout the entire mission. Classical CBFs are often applied via quadratic programs (QPs) to enforce safety while minimally adjusting nominal control, enabling real-time collision avoidance for robots, manipulators, and UAVs. Similarly, the existence of higher-order CBF (HOCBF) extends CBFs to handle higher relative-degree systems [34], which is crucial for UAVs with second-order dynamics. The VOCBF [13] encodes safety in velocity space. Moreover, the CBF is integrated with differentiable optimization-based methods to enable smoother control pipelines [6]. Similarly, a probabilistic enumeration is introduced for stochastic risk allocation [19]. Runtime adaptive safety filtering and robust formulations have further broadened the applicability of CBFs in robotics [14, 7].
@@
Despite this progress, current CBF implementations face limitations for DOA in UAVs. Existing CBF formulations act at the position or velocity level [13, 23], but do not provide closed-form acceleration-space constraints compatible with fast low-level thrust control. The HOCBFs do not directly address acceleration or thrust for UAV DOA [28]. Although their constraints are solver-independent, online QP approaches can limit real-time use on embedded platforms [34]. The current CBF is insufficient to address latency and uncertainty in UAV DOA. The ISSf-CBF [1] and tube-based formulations [29] consider robustness to input delay, but they are conservative and rely on iterative solvers. Probabilistic CBFs improve risk handling, but have not been adapted to acceleration-level constraints or integrated with hard real-time systems [19, 14]. Similarly, CBF with chance-constraints emphasizes probabilistic risk margins for UAVs but remains velocity-centric [36]. The QP-based CBF solvers and sampling-based planners are too slow for high-rate DOA control, particularly on resource-constrained UAV hardware [23, 7, 34]. Lightweight and computationally efficient MPC extensions, such as TinyMPC, proving formal guarantees for fast reaction, are available for Crazyflies, but they do not quantify performance in high closing velocity DOA scenarios [25].
@@
In this paper, we address the current CBF-QP limitation for DOA in UAVs with a distributionally robust acceleration control barrier function (DR-ACBF). DR-ACBF directly restricts UAV acceleration commands through safe, linear, half-space constraints. The half-space constraints specify the safe side of the projection plane for the UAV’s acceleration, ensuring that the UAV remains within this region. We demonstrate that enforcing a linear acceleration-space constraint derived from the second-order CBF dynamics satisfies the original CBF inequality over a short horizon under bounded uncertainty and latency assumptions. Additionally, the perception-to-actuation latency and uncertainty are handled by incorporating clearance buffering [9] and probabilistic tightening based on Cantelli inequalities [27], similar in spirit to stochastic CBFs [19, 14] but specialized for acceleration space. Moreover, an early-warning trigger layer based on distributionally robust conditional value at risk (DR-CVaR) [24, 31, 5] anticipates risk growth within the same maximum horizon $H_{max}$ and inflates safety margins before any violation becomes imminent. This trigger improves the robustness of ACBF by ensuring that subsequent ACBF enforcement remains feasible and effective under latency- and jerk-limited execution. We use fixed-time Gauss-Southwell projection [26, 12] instead of online QPs, enabling real-time control at a rate of 100 Hz.
@@
In contrast to existing CBF-driven collision avoidance frameworks, our DR-ACBF framework provides: (i) explicit acceleration-space tightening to stay safe, (ii) low computational cost via iterative projection, and (iii) probabilistic and latency-aware safety margins. We derive a closed-form line-of-sight (LoS) acceleration half-space with horizon-matched gains and a fixed-time projection enforcement with Gauss-Southwell. To the best of our knowledge, no prior works present an acceleration-space CBF filter with latency and risk tightening for UAV obstacle avoidance. The DR-ACBF achieves real-time, reliable UAV obstacle avoidance via acceleration-level CBFs, DR-CVaR trigger, and fixed-time projection, reducing command time by over 32% while maintaining high safety and success rates as evaluated in a simulation study. Projection-based DR-ACBF reduces processing time by 31% as compared to QP-based solvers. Experiments with Crazyflie UAVs confirm the feasibility of our approach.
@@ II Problem Formulation
Figure 1: Illustration of the proposed DR-ACBF framework. The UAV avoids dynamic obstacles (top) by constraining its acceleration within distributionally robust half-spaces marked in grey. The UAV view (bottom) depicts the projected half-space plane towards the obstacles in FoV. The effective clearance $R_{\mathrm{eff}}$ accounts for latency and actuation limits, while the obstacles’ uncertainty ellipsoids $\sigma_{dH,i}$ and risk levels $\alpha_{i}$ tighten the constraints under sensing uncertainty. The safe set of action $C$ ensures that the UAV’s safe zone is not violated.
@@ II-A Problem Statement
We model the UAV dynamics by second-order differential equations for simpler control design. We consider an acceleration-controlled quadrotor UAV operating in an environment with $N$ dynamic obstacles. The obstacles in the sensing field of view (FoV) are detected and tracked with typical trackers (e.g., SAFE-IMM tracker [18]), which deliver state estimates of the obstacles. The input space consists of the position $p_{e}\in\mathbb{R}^{3}$, velocity $v_{e}\in\mathbb{R}^{3}$, and acceleration $a_{e}\in\mathbb{R}^{3}$ of the UAV as well as the relative position, velocity, and acceleration of the obstacle. A nominal controller provides an acceleration command $a_{\mathrm{n}}\in\mathbb{R}^{3}$, while the UAV is subject to perception-to-actuation latency $\Delta\!\geq\!0$ and bounded actuation, i.e., limits for acceleration $a_{\max}\!>\!0$, jerk $j_{\max}\!>\!0$, and speed $v_{\max}\!>\!0$. The objective of DR-ACBF is to design a real-time safety filter that modifies $a_{\mathrm{n}}$ to a safe command $a_{e}\in\mathbb{R}^{3}$ while enforcing probabilistic collision avoidance under a user-specified global risk budget $\alpha_{\mathrm{total}}$ (see Figure 1).
@@ II-D Classical CBF Formulation
Ames et al. [2, 3] introduced the classical CBF-QP architecture, which guarantees that an autonomous agent moves toward its goal while maintaining safety constraints. Let $x\in\mathbb{R}^{n}$ denote the system state and $u\in\mathbb{R}^{m}$ the control input. The control-affine dynamics are
@@
When (9) is satisfied (i.e., the CBF constraint is enforced), the safe set $\mathcal{C}$ is forward invariant, meaning $x(0)\in\mathcal{C}$ implies $x(t)\in\mathcal{C}$ for all $t\geq 0$. At each control step, one computes $u^{\star}$ by solving the QP
$u^{\star}=\arg\min_{u}\tfrac{1}{2}\|u-u_{\mathrm{nom}}\|^{2}$ (10)
$\text{s.t.}\quad L_{f}h(x)+L_{g}h(x)u+\alpha(h(x))\geq 0$
where $u_{\mathrm{nom}}$ is the nominal control. The CBF inequality (9) strictly enforces the safety required to reach the goal successfully. We omit the control Lyapunov function, as no soft performance constraint is used.
@@ III Method
Figure 2: System block diagram. The process begins with sensing the UAV and obstacle states. UAV control generates a nominal acceleration command $a_{n}$, which is corrected by DR-ACBF avoidance. The DR-CVaR layer provides an early warning and requires acceleration estimation and risk allocation to compute the safety constraints. Finally, ACBF and projection replace $a_{n}$ with the safe acceleration command $a_{e}$ to avoid obstacles.
@@ III-A Acceleration CBF
We model the UAV dynamics similar to Equation (1) with $\|a_{e}\|\leq a_{\max}$:
@@
The combined sum of obstacle and UAV radius is $R_{\mathrm{sum}}=R_{\mathrm{uav}}+R_{\mathrm{obs}}$ as illustrated in Figure 1. The barrier function for distance is then
@@
where $k_{1}$ damps the approach velocity and $k_{2}$ penalizes distance violation. In order to project the second-order condition into an acceleration half-space, we linearize it over a short lookahead horizon $H>0$. Where the $a_{i}$ is not modeled; its worst-case effect is padded into $R_{\mathrm{eff},i}$ and $v_{i}$ enters via $v_{i}^{\mathrm{cl}}$, yielding a conservative, closed-form affine constraint enforceable at control rate without a QP in the UAV acceleration
@@
where $R_{\mathrm{eff},i}$ is the effective clearance that accounts for sensing and actuation latencies. Each obstacle thus defines one linear half-space of safe accelerations. Normalization by $\|P_{i}\|$ is already included in $n_{i}$, so (16) requires no extra factor.
The effective clearance in distance is enlarged to account for perception latency, actuation limits, and jerk bound
@@
where $v_{\max}$ is the UAV speed limit, $a_{\max}$ and $j_{\max}$ are acceleration and jerk bounds, $\Delta$ is total perception-to-actuation latency, and the $\hat{a}_{\max,i}$ is the learned bound on obstacle acceleration. The half-space linearization assumes nearly constant LoS direction and bounded constant acceleration over $H$, and the neglected higher-order effects are conservatively covered by inflating $R_{\mathrm{sum}}$ to $R_{\mathrm{eff}}$ (especially for small $H$). The sliding-mode differentiator (SMD) is used to compute the $\hat{a}_{\max,i}$ with a user-defined $a_{\max,i}$, aiming to keep it large to address the worst possible acceleration of obstacles. This is explained in detail in the next sections.
To compensate for latency variations, the latency term in $R_{\mathrm{eff}}$ is bounded by $v_{\max}\Delta+\tfrac{1}{2}a_{\max}\Delta^{2}$. While the new command is held over $H$ with bounded jerk $j_{\max}$, the worst additional shrink from the acceleration ramp is upper-bounded by $\tfrac{1}{6}j_{\max}H^{3}$. The obstacle displacement over $H$ is limited by $\tfrac{1}{2}\hat{a}_{\max,i}H^{2}$. Summing up all these components to $R_{\mathrm{sum}}$ yields (17). The $\Delta$ includes pre-actuation latency to account for delays in activating the controls. The impact of such latency on safe high-speed navigation with bounded acceleration is analyzed in [9]. The trackers provide estimates of position and velocity [18]. Therefore, if we consider the position and velocity covariance from the tracker, the LoS variance is bounded by Eq. (6). Over a short lookahead horizon $H$, the coefficients $k_{1}=2/H$ and $k_{2}=2/H^{2}$ ensure that any acceleration satisfying the affine constraint maintains a non-decreasing barrier value of $h_{i}(t)$ over $H$. Using the 2nd order Taylor approximation $h(t+H)\approx h(t)+H\dot{h}(t)+\tfrac{1}{2}H^{2}\ddot{h}(t)$, enforcing $h(t+H)\geq 0$ yields $\ddot{h}(t)+\tfrac{2}{H}\dot{h}(t)+\tfrac{2}{H^{2}}h(t)\geq 0$, hence $k_{1}=\tfrac{2}{H}$ and $k_{2}=\tfrac{2}{H^{2}}$.